# Tuning the Frequencies: Robust Training for Sinusoidal Neural Networks

### v2

@@ 6 Conclusions and limitations
We presented a study based on Fourier theory for sinusoidal INRs and used it to propose new schemes for both initialization and spectral control. Positively, our theoretical contributions generalize to different domains and aim to provide a better understanding of sinusoidal INRs for all applications. While we limited our experiments to images to demonstrate our theoretical claims, future works include using our schemes for other signals. We also have focused our studies on a single hidden layer architecture leaving the investigation of deep architectures for a follow-up work. In that sense, we are certainly considering future research aiming to extrapolate our results to general architectures and domains. Thus, we consider this work as a first step to unravel the powerful expressiveness of sinusoidal INRs.

### v3

@@ 6 Conclusions and limitations
We presented a study based on Fourier theory for sinusoidal MLPs, which resulted in novel, robust schemes for both initialization and spectral control of such networks. While our experiments focused on images to validate our theoretical claims, future work includes applying our schemes to other signal types. Additionally, we have focused our studies on 3-layer MLPs, leaving the investigation of deep networks for future research. We consider this work a first step toward unraveling the powerful expressiveness of sinusoidal MLPs.
Acknowledgments
We would like to thank CAPES, FAPERJ, and Google for partially funding this work.